# mDP3: A Training-free Approach for List-wise Frame Selection in Video-LLMs

### v1

@@ Abstract
Video large language models (Video-LLMs) have made significant progress in understanding videos. However, processing multiple frames leads to lengthy visual token sequences, presenting challenges such as the limited context length cannot accommodate the entire video, and the inclusion of irrelevant frames hinders visual perception. Hence, effective frame selection is crucial. This paper emphasizes that frame selection should follow three key principles: query relevance, list-wise diversity, and sequentiality. Existing methods, such as uniform frame sampling and query-frame matching, do not capture all of these principles. Thus, we propose Markov decision determinantal point process with dynamic programming (mDP3) for frame selection, a training-free and model-agnostic method that can be seamlessly integrated into existing Video-LLMs. Our method first estimates frame similarities conditioned on the query using a conditional Gaussian kernel within the reproducing kernel Hilbert space (RKHS). We then apply the determinantal point process (DPP) to the similarity matrix to capture both query relevance and list-wise diversity. To incorporate sequentiality, we segment the video and apply DPP within each segment, conditioned on the preceding segment selection, modeled as a Markov decision process (MDP) for allocating selection sizes across segments. Theoretically, mDP3 provides a $(1-1/e)$-approximate solution to the NP-hard list-wise frame selection problem with pseudo-polynomial time complexity, demonstrating its efficiency. Empirically, mDP3 significantly outperforms existing methods, verifying its effectiveness and robustness.
@@ 1 Introduction
Building on the insights from the above analysis, we emphasize that effective frame selection in Video-LLMs should adhere to three key principles: query relevance, list-wise diversity, and sequentiality. Previous research has addressed some of these principles but has not fully encompassed all of them, particularly the latter two [16, 54]. Modeling list-wise relationships in subset frame selection presents two main challenges: 1) the lack of available supervised list-wise datasets; 2) even if a dataset is constructed to learn a list-wise scoring function, finding the optimal subset with the highest score remains an NP-hard problem that cannot be solved in polynomial time complexity [4, 6].
In this paper, we propose a training-free, model-agnostic list-wise frame selection method called Markov Decision Determinantal Point Process with Dynamic Programming, abbreviated as mDP3. We reuse pretrained VLMs to compute list-wise score for the subset using Determinantal Point Process (DPP), thereby eliminating the need for collecting a supervised dataset or training a scoring model from scratch. DPP was originally introduced to characterize the Pauli exclusion principle, where no two fermions can occupy the same quantum state. This repulsion, interpreted as list-wise diversity, is effectively modeled by DPP [37], which can obtain a $(1-1/e)$-approximate solution with polynomial complexity [6]. In mDP3, we extend standard DPP to consider query relevance, list-wise diversity, and sequentiality for frame selection in VidQA.
@@
To evaluate mDP3, we integrate it as a plug-and-play process during the inference stage of state-of-the-art (SOTA) Video-LLMs, including VILA-V1.5 [28], MiniCPM-V2.6 [63], and LLaVA-OneVision [21]. We conduct extensive experiments on three widely used long-video benchmarks: Video-MME [13], MLVU [71], and LongVideoBench [58]. Experimental results show significant performance improvements over baseline methods.
@@
    mDP3 is effective, offering a $(1-1/e)$-approximate solution to the NP-hard list-wise subset frame selection problem, achieving significant performance improvements over existing baselines.
@@
    mDP3 is training-free and model-agnostic, enabling seamless plug-and-play integration with various existing Video-LLMs.
@@ 2 Realted Work
Most existing Video-LLMs rely on uniform frame sampling [8, 23, 33, 46], which arbitrarily disregards the current query, often missing relevant frames and including irrelevant ones. Traditional key-frame methods that detect significant transitions between adjacent frames also face this limitation [45, 40, 57]. Some studies have attempted to use the query to retrieve the top-$k$ most relevant frames with VLMs but fail to consider frame diversity [51, 65], often resulting in redundant frame selection, and increased computational burden without meaningful information gain. Some cluster-based methods attempt to regularize pair-wise diversity among frames but still fail to capture combinatorial relationships in a list-wise manner [17, 16]. Recently, Yu et al. [66] proposed a frame selection method tailored for VILA [28], which requires collecting a supervised dataset and end-to-end training of a selector, limiting its generalizability to other Video-LLMs. Furthermore, this method does not explicitly address list-wise diversity or sequentiality, both of which are essential for effective frame selection.
@@ 3.1 High-dimensional Conditional Similarity
Instead of directly estimating similarity using cosine similarity or the inner product of embeddings, we map the embeddings into an RKHS via the reproducing kernel feature map $\phi\in\mathcal{H}_{k}$, where, according to the kernel trick [44], the high-dimensional similarity in RKHS is expressed by $\phi(x),\phi(y)\right>=k(x,y)$. We employ a characteristic multi-kernel defined as a convex combination of multiple positive semi-definite (PSD) kernels $\{k_{u}\}$ (e.g. Gaussian kernel):
@@ 3.2 Determinantal Point Process (DPP)
Frame selection follows the law of diminishing marginal utility, as the marginal gain from selecting additional frames is non-increasing (as shown in Fig. 1). This phenomenon is formally known in statistics as submodularity, which is a well-established property of DPP, as shown in previous research [14, 18]. Selecting the optimal subset to maximize the submodular score function Eqs. 7 and 9, known as the submodular maximization problem in subset selection, is generally NP-hard. Fortunately, there is a popular greedy algorithm can solve this problem, which offers polynomial complexity and guarantees a $(1-1/e)$-approximation, i.e., the solution is at least $(1-1/e)$ of the optimal score [6].
@@ 3.3 Markov Decision DPP
First, we segment a video into multiple segments of equal length $m$, so that a video with $n$ frames is divided into $T=\lceil n/m\rceil$ segments. The corresponding frame indices of the video, $\mathbb{N}_{n}$, are divided as $\mathbb{N}_{n}=\bigcup_{t=1}^{T}N_{t}$, where $N_{t}$ denotes the frame indices of the $t$-th segment, with $|N_{t}|=m$. Moreover, let the selected subset from the $t$-th segment be denoted as $S_{t}\subseteq N_{t}$. Therefore, when applying DPP to the $t$-th segment, conditioned on the selection in the $(t-1)$-th segment $S_{t-1}$, the DPP conditional distribution [15] for the current segment $S_{t}$ is expressed as:
@@
Determining the optimal allocation of selection size $k_{t}$ for each segment is challenging. We observe that the MAP inference (Eq. 13) of the DPP conditional distribution (Eq. 11) uniquely depends on a triplet state, consisting of: 1) the candidate indices in the current segment $N_{t}$, 2) the selections from the preceding segment $S_{t-1}^{*}$, determined by $k_{t-1}$, and 3) the remaining capacity for further selections $k-C_{t-1}$, where $C_{t-1}$ denotes the total selection size before the $(t-1)$-th (inclusive) segment. Consequently, this MAP inference can be formulated as an MDP, where the state is defined as the triplet $(N_{t},k_{t-1},C_{t-1})$. Moreover, the segment index $t-1$ can uniquely determine $N_{t}$. Therefore, the state can be represented as $(t\in[1,T],k_{t}\in[0,k],C_{t}\in[0,k])$. The state transition from $t-1$ to $t$ is deterministic when the action of selecting $k_{t}$ frames from the $t$-th segment is given, which can be expressed as:
$(t-1,k_{t-1},C_{t-1})\xrightarrow{\scriptsize\begin{array}[]{c}\text{selecting}S_{t}\\ \text{with }|S_{t}|=k_{t}\end{array}}(t,k_{t},\overbrace{C_{t-1}+k_{t}}^{C_{t}})\,.$ (14)
@@
Using this definition and the $(1-1/e)$-approximate greedy DPP algorithm, we can apply a dynamic programming approach to identify a solution $\hat{S}$ such that $\mathrm{Score}(\hat{S})\geq(1-1/e)\cdot\mathrm{Score}(S^{*})$, with a closely $\mathcal{O}(nk^{4})$ time complexity (proof provided in the Appendix).
@@ 17                         Update $\mathcal{T}_{t,C_{t}}\leftarrow\mathcal{T}_{t-1,C_{t-1}}\cup S_{t}$;
Model LLM Size #Frames Video-MME${}_{(wo/w\ subs)}$ MLVU LVBval
@@
Dynamic programming can be used to update the $Q$ score and identify the solution $\hat{S}$ with an optimal allocation of $k$, with a time complexity closely $\mathcal{O}(nk^{4})$. In practice, to improve efficiency further, we adopt lazy dynamic programming strategies. First, we directly maintain the $Q^{*}$ score instead of updating $Q$. This strategy assumes that selecting a frame from the $t$-th segment depends only on the preceding state with optimal $k_{t-1}^{*}$, i.e., $(t-1,k_{t-1}^{*},C_{t-1})$, corresponding to $Q^{*}(t-1,C_{t-1})$, similar to the online MAP inference described in Eq. 13 [15]. Second, only the most recently selected frame is used as the condition in Eq. 15, approximating an adaptive condition size and a fixed candidate segment size $m$ for the $t$-th segment. Consequently, with the lazy strategies, the state is simplified to $(t,C_{t})$. The policy function is denoted by $S_{t}^{*}=\pi(t,C_{t})$, and $\mathcal{T}_{t,C_{t}}$ is an initially empty list [][\,][ ] that tracks the selection trace $[S_{1},S_{2},\dots,S_{t}]$. The corresponding reward simplifies to:
@@
The pseudocode for mDP3 is provided in Algo. 1. The end index $(t,C_{t}+k_{t})$ is iterated, while the start index $(t-1,C_{t})$ remains fixed. The update functions are equivalent with Sec. 3.3. This implementation is more efficient as the greedy DPP algorithm (line 9) can be computed incrementally in $\mathcal{O}(mk)$. Thus, the worst-case time complexity of Algo. 1 is closely $\mathcal{O}(nk^{3})$ in practice (proof provided in Appendix). Compared to the standard DPP complexity of $\mathcal{O}(nk^{2})$, the additional factor $k\ll n$ is negligible. Additionally, the iteration in line 6 is independent, enabling parallel updates. This results in a time complexity of $\mathcal{O}(nk^{2})$, making its efficiency comparable to the standard DPP.
@@ 4 Experiments
We integrate mDP3 as a plug-and-play process during inference with 7B SOTA Video-LLMs. We conduct experiments using LMMs-Eval [20] and VLMEvalKit [11] on three long video benchmarks. Due to space limitations, additional experimental details and results are provided in the Appendix.
@@
VILA-V1.5 47.1+3.7$\uparrow$ 56.1+1.6$\uparrow$ 60.5+5.8$\uparrow$ 43.4+4.9$\uparrow$ 42.7+2.9$\uparrow$
+mDP3 50.8+3.7$\uparrow$ 57.7+1.6$\uparrow$ 66.3+5.8$\uparrow$ 48.3+4.9$\uparrow$ 45.6+2.9$\uparrow$
@@ 4.1 Experiments Protocol
Backbone Models & Primary Baselines: Based on the Video-MME leaderboard [13], we choose the 7B SOTA Video-LLMs LLaVA-OneVision-7B [21] and MiniCPM-V2.6-7B [63] as our primary baselines. To compare with the recent frame-selection approach [66], tailor-made for VILA [28], we also include VILA-V1.5-7B as the primary baseline. We integrate mDP3 with these baselines and evaluate its performance against other recent Video-LLMs.
@@
Implementations Details: Following the setup in Yu et al. [66], we uniformly sample 128 candidate frames and apply mDP3 to select 8 frames as input to the Video-LLMs. All three primary baselines employ SigLIP [67] as the visual encoder, which we also adopt as the VLM in Eq. 2 to avoid including excessive additional parameters. Since the context length of the text encoder in SigLIP is 64, potentially insufficient for the entire question, we split the text sequence into multiple sequences of equal length, each no longer than 64 tokens. We then extract multiple text embeddings and aggregate them into a final text embedding using pooling. Additionally, we set the trade-off hyperparameter $\lambda=0.2$ and the segment size $m=32$ for all tasks and benchmarks in mDP3, determined through cross-validation on a subset of the LLaVA-OneVision-mid [21] training set.
@@ 4.2.1 Comparison with SOTA Methods
Tab. 1 presents a comparison of mDP3 while applied to VILA-V1.5 [28], MiniCPM-V2.6 [63], and LLaVA-OneVision [21]. Additionally, we compare mDP3 to recent Video-LLMs, including Video-LLaVA [27], Qwen-VL-Chat [3], VideoChat2 [23], Chat-UniVi-V1.5 [16], VideoLLaMA2 [9], LLaVA-NEXT-QW2 [30], LongVILA [61], LongVA [68], and Video-XL [46]. The number of input frames for the LLMs is shown in Tab. 1. Specifically, Video-XL processes 128 frames for Video-MME and 256 frames for MLVU, while LLaVA-OneVision reports the official results using a well-tuned frame count. Compared to the baselines, mDP3 substantially improves the SOTA performance across all three evaluated benchmarks. The main results presented in Tab. 1 highlight several key observations:
@@
    Our frame selection method, mDP3, empowers LLaVA-OneVision to achieve the highest overall performance without subtitles, while MiniCPM-V2.6 excels with subtitles, confirming the effectiveness of mDP3.
@@
    When mDP3 is applied to Video-LLMs, including VILA, MiniCPM-V2.6, and LLaVA-OneVision, it consistently outperforms uniform frame sampling. This demonstrates that mDP3 is effective in frame selection and generalizes well across various Video-LLMs.

### v2

@@ Abstract
Video large language models (Video-LLMs) have made significant progress in understanding videos. However, processing multiple frames leads to lengthy visual token sequences, presenting challenges such as the limited context length cannot accommodate the entire video, and the inclusion of irrelevant frames hinders visual perception. Hence, effective frame selection is crucial. This paper emphasizes that frame selection should follow three key principles: query relevance, list-wise diversity, and sequentiality. Existing methods, such as uniform frame sampling and query-frame matching, do not capture all of these principles. Thus, we propose Markov decision determinantal point process with dynamic programming (mDP3) for frame selection, a training-free and model-agnostic method that can be seamlessly integrated into existing Video-LLMs at test time. Our method first estimates frame similarities conditioned on the query using a conditional Gaussian kernel within the reproducing kernel Hilbert space (RKHS). We then apply the determinantal point process (DPP) to the similarity matrix to capture both query relevance and list-wise diversity. To incorporate sequentiality, we segment the video and apply DPP within each segment, conditioned on the preceding segment selection, modeled as a Markov decision process (MDP) for allocating selection sizes across segments. Theoretically, mDP3 provides a (11/e)(1-1/e)-approximate solution to the NP-hard list-wise frame selection problem with pseudo-polynomial time complexity, demonstrating its efficiency. Empirically, mDP3 significantly outperforms existing methods, verifying its effectiveness and robustness.
@@ 1 Introduction
Building on the insights from the above analysis, we emphasize that effective frame selection in Video-LLMs should adhere to three key principles: query relevance, list-wise diversity, and sequentiality. Prior research has addressed some of them but has not fully encompassed all of them, particularly the latter two [17, 57, 51, 16]. Modeling list-wise relationships in subset frame selection presents two main challenges: 1) the lack of available supervised list-wise datasets; 2) even if a dataset is constructed to learn a list-wise scoring function, finding the optimal subset with the highest score remains an NP-hard problem that cannot be solved in polynomial time complexity [4, 6].
In this paper, we propose a training-free, model-agnostic list-wise frame selection method called Markov Decision Determinantal Point Process with Dynamic Programming, abbreviated as mDP3. We reuse pretrained VLMs to compute list-wise score for the subset using Determinantal Point Process (DPP), thereby eliminating the need for collecting a supervised dataset or training a scoring model from scratch. DPP was originally introduced to characterize the Pauli exclusion principle, where no two fermions can occupy the same quantum state. This repulsion, interpreted as list-wise diversity, is effectively modeled by DPP [38], which can obtain a (11/e)(1-1/e)-approximate solution with polynomial complexity [6]. In mDP3, we extend standard DPP to consider query relevance, list-wise diversity, and sequentiality for frame selection in VidQA.
@@
To evaluate mDP3, we integrate it as a plug-and-play process at the test time of state-of-the-art (SOTA) Video-LLMs, including VILA-V1.5 [29], MiniCPM-V2.6 [66], LLaVA-OneVision [22], and Ovis2 [52]. We conduct extensive experiments on three widely used long-video benchmarks: Video-MME [13], MLVU [75], and LongVideoBench [61]. Experimental results show significant performance improvements over baseline methods.
@@
    mDP3 is effective, offering a (11/e)(1-1/e)-approximate solution to the NP-hard list-wise subset frame selection problem, achieving significant performance improvements over existing baselines.
@@
    mDP3 is training-free and model-agnostic, enabling seamless plug-and-play integration with various existing Video-LLMs at test time.
@@ 2 Realted Work
Most existing Video-LLMs rely on uniform frame sampling [8, 24, 34, 47], which arbitrarily disregards the current query, often missing relevant frames and including irrelevant ones. Traditional key-frame methods that detect significant transitions between adjacent frames also face this limitation [46, 41, 60]. Some studies have attempted to use the query to retrieve the top-$k$ most relevant frames with VLMs but fail to consider frame diversity [54, 69], often resulting in redundant frame selection, and increased computational burden without meaningful information gain. Some cluster-based methods attempt to regularize pair-wise diversity among frames but still fail to capture combinatorial relationships in a list-wise manner [18, 17]. Recently, Yu et al. [70] proposed a frame selection method tailored for VILA [29], which requires collecting a supervised dataset and end-to-end training of a selector, limiting its generalizability to other Video-LLMs. Furthermore, these method does not explicitly address list-wise diversity or sequentiality, both of which are essential for effective frame selection.
@@ 3.1 High-dimensional Conditional Similarity
Instead of directly estimating similarity using cosine similarity or the inner product of embeddings, we map the embeddings into an RKHS via the reproducing kernel feature map $\phi\in\mathcal{H}_{k}$, where, according to the kernel trick [45], the high-dimensional similarity in RKHS is expressed by $\left<\phi(x),\phi(y)\right>=k(x,y)$. We employ a characteristic multi-kernel defined as a convex combination of multiple positive semi-definite (PSD) kernels $\{k_{u}\}$ (e.g. Gaussian kernel):
@@ 3.2 Determinantal Point Process (DPP)
Frame selection follows the law of diminishing marginal utility, as the marginal gain from selecting additional frames is non-increasing (as shown in Fig. 1). This phenomenon is formally known in statistics as submodularity, which is a well-established property of DPP, as shown in previous research [14, 19]. Selecting the optimal subset to maximize the submodular score function Eqs. 7 and 9, known as the submodular maximization problem in subset selection, is generally NP-hard. Fortunately, there is a popular greedy algorithm can solve this problem, which offers polynomial complexity and guarantees a (11/e)(1-1/e)-approximation, i.e., the solution is at least (11/e)(1-1/e) of the optimal score [6].
@@ 3.3 Markov Decision DPP
First, we segment a video into multiple segments of equal length $m$, so that a video with $n$ frames is divided into $T=\lceil n/m\rceil$ segments. The corresponding frame indices of the video, $\mathbb{N}_{n}$, are divided as $\mathbb{N}_{n}=\bigcup_{t=1}^{T}N_{t}$, where $N_{t}$ denotes the frame indices of the $t$-th segment, with $|N_{t}|=m$. Moreover, let the selected subset from the $t$-th segment be denoted as $S_{t}\subseteq N_{t}$. Therefore, when applying DPP to the $t$-th segment, conditioned on the selection in the (t1)(t-1)-th segment $S_{t-1}$, the DPP conditional distribution [15] for the current segment $S_{t}$ is expressed as:
@@
Determining the optimal allocation of selection size $k_{t}$ for each segment is challenging. We observe that the MAP inference (Eq. 13) of the DPP conditional distribution (Eq. 11) uniquely depends on a triplet state, consisting of: 1) the candidate indices in the current segment $N_{t}$, 2) the selections from the preceding segment $S_{t-1}^{*}$, determined by $k_{t-1}$, and 3) the remaining capacity for further selections $k-C_{t-1}$, where $C_{t-1}$ denotes the total selection size before the (t1)(t-1)-th (inclusive) segment. Consequently, this MAP inference can be formulated as an MDP, where the state is defined as the triplet $(N_{t},k_{t-1},C_{t-1})$. Moreover, the segment index t1t-1 can uniquely determine $N_{t}$. Therefore, the state can be represented as $(t\in[1,T],k_{t}\in[0,k],C_{t}\in[0,k])$. The state transition from t1t-1 to $t$ is deterministic when the action of selecting $k_{t}$ frames from the $t$-th segment is given, which can be expressed as:
$(t-1,k_{t-1},C_{t-1})\xrightarrow{\hskip-2.84544pt\scriptsize\begin{array}[]{c}\text{selecting }S_{t}\\ \text{with }|S_{t}|=k_{t}\end{array}\hskip-2.84544pt}(t,k_{t},\overbrace{C_{t-1}+k_{t}}^{C_{t}})\,.$ (14)
@@
Using this definition and the (11/e)(1-1/e)-approximate greedy DPP algorithm, we can apply a dynamic programming approach to identify a solution $\hat{S}$ such that $\mathrm{Score}(\hat{S})\geq(1-1/e)\cdot\mathrm{Score}(S^{*})$, with a closely $\mathcal{O}(nk^{4})$ time complexity (proof provided in the Appendix).
@@ 17            Update $\mathcal{T}_{t,C_{t}}\leftarrow\mathcal{T}_{t-1,C_{t-1}}\cup S_{t}$;
Model
LLM
Size
[…]
@@
Ovis2 7B 8 58.9 / 62.1 67.8 / 72.0 57.4 /60.3 51.3 / 54.0 60.9 56.9
[…]
Table 1: Comparison of Video-LLMs with and without mDP3. LVBval denotes the LongVideoBench validation set without subtitle interleaving. Video-XL uses 128/256 frames on Video-MME/MLVU. LLaVA-OneVision refers to official report with well-tuned frames.
Dynamic programming can be used to update the $Q$ score and identify the solution $\hat{S}$ with an optimal allocation of $k$, with a time complexity closely $\mathcal{O}(nk^{4})$. In practice, to improve efficiency further, we adopt lazy dynamic programming strategies. First, we directly maintain the $Q^{*}$ score instead of updating $Q$. This strategy assumes that selecting a frame from the $t$-th segment depends only on the preceding state with optimal $k_{t-1}^{*}$, i.e., $(t-1,k_{t-1}^{*},C_{t-1})$, corresponding to $Q^{*}(t-1,C_{t-1})$, similar to the online MAP inference described in Eq. 13 [15]. Second, only the most recently selected frame is used as the condition in Eq. 15, approximating an adaptive condition size and a fixed candidate segment size $m$ for the $t$-th segment. Consequently, with the lazy strategies, the state is simplified to $(t,C_{t})$. The policy function is denoted by $S_{t}^{*}=\pi(t,C_{t})$, and $\mathcal{T}_{t,C_{t}}$ is an initially empty list [][\,] that tracks the selection trace $S_{1},S_{2},\dots,S_{t}]$. The corresponding reward simplifies to:
@@
The pseudocode for mDP3 is provided in Algorithm 1. The end index $(t,C_{t}+k_{t})$ is iterated, while the start index $(t-1,C_{t})$ remains fixed. The update functions are equivalent with Sec. 3.3. This implementation is more efficient as the greedy DPP algorithm (line 9) can be computed incrementally in $\mathcal{O}(mk)$. Thus, the worst-case time complexity of Algorithm 1 is closely $\mathcal{O}(nk^{3})$ in practice (proof in Appendix). Compared to the standard DPP complexity of $\mathcal{O}(nk^{2})$, the additional factor $k\ll n$ is negligible. Additionally, the iteration in line 6 is independent, enabling parallel updates. This results in a time complexity of $\mathcal{O}(nk^{2})$, making its efficiency comparable to the standard DPP.
@@ 4 Experiments
We integrate mDP3 as a plug-and-play process during inference with 7B SOTA Video-LLMs. Following the setup in Yu et al. [70], we uniformly sample 128 candidate frames and apply mDP3 to select 8 frames as input to the Video-LLMs. We conduct experiments using LMMs-Eval [21] and VLMEvalKit [11] on three long video benchmarks. Due to space limitations, additional experimental details and results are provided in the Appendix. Code in https://github.com/sunh-23/MDP3.
@@ 4.1 Experiments Protocol
Backbone Models & Primary Baselines: Based on the Video-MME leaderboard [13], we choose the 7B SOTA Video-LLMs LLaVA-OneVision-7B [22], MiniCPM-V2.6-7B [66] and Ovis2-8B [52] as our primary baselines. To compare with the recent frame selection approach [70], tailor-made for VILA [29], we also include VILA-V1.5-7B as the primary baseline. We integrate mDP3 with these baselines and evaluate it against other recent Video-LLMs.
@@ 4.2.1 Comparison with SOTA Methods
Tab. 1 presents a comparison of mDP3 while applied to VILA-V1.5 [29], MiniCPM-V2.6 [66], LLaVA-OneVision [22] and Ovis2 [52]. We compare mDP3 to recent Video-LLMs: Video-LLaVA [28], Qwen-VL-Chat [3], VideoChat2 [24], Chat-UniVi-V1.5 [17], VideoLLaMA2 [9], LLaVA-NEXT-QW2 [31], LongVILA [64], LongVA [72], and Video-XL [47]. The number of input frames for the LLMs is shown in Tab. 1. Specifically, Video-XL processes 128 frames for Video-MME and 256 frames for MLVU, while LLaVA-OneVision reports the official results using a well-tuned frame count. Compared to the baselines, mDP3 substantially improves the SOTA performance across all three evaluated benchmarks. The main results presented in Tab. 1 highlight several key observations:
@@
    Our frame selection method, mDP3, empowers Ovis2 to achieve the highest overall performance across various benchmarks, confirming the effectiveness of mDP3.
@@
    When mDP3 is applied to the various primary Video-LLMs, it consistently outperforms uniform frame sampling. This validates that mDP3 is effective in frame selection and generalizes well across various Video-LLMs.